# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v1

@@ 8 Discussion
A longstanding view in computational and cognitive neuroscience treats perception as predictive and time-coupled: slowly varying, context-invariant structure coexists with fast-changing, surprising residuals (Levy, 2008; Millidge et al., 2024; Shain et al., 2020). We find that this structure is reflected in LM activations (Sec. 3): activations are strongly correlated with recent context and can be decomposed accordingly. In contrast, SAEs assume independence over time; this is misaligned with the identified non-stationary structure in language model activations. Our protocol—Temporal Feature Analysis—relaxes this assumption, incorporating empirically observed temporal correlations to surface structure that i.i.d. priors obscure. We showed that Temporal Feature Analysis is able to highlight temporal structures in its predictable codes that SAEs are unable to capture due to their prior-data mismatch. Overall, our work adds to recent findings that inductive biases about the specific application under study are necessary to build improved interpretability tools (Hindupur et al., 2025; Costa et al., 2025; Sutter et al., 2025).
@@ 8.1 Future Work
We believe our work brings interpretability of language models one step closer to recent advances in interpretability of neural networks designed to simulate computational neuroscience tasks (Sussillo and Barak, 2013; Cotler et al., 2023; Mante et al., 2013; Maheswaranathan et al., 2019; Maheswaranathan and Sussillo, 2020). Specifically, rather than isolated directions, as is assumed in both sparse coding (Olshausen and Field, 1996; 1997) and SAEs (Elhage et al., 2022; Bricken et al., 2023), “features” are better viewed as low-dimensional manifolds that evolve over time—as has also been stated in other recent work (Modell et al., 2025; Fel et al., 2025; Gurnee et al., 2025). In particular, we see in our results that predictive codes trace smooth, stable trajectories (events, roles, long-horizon constraints), whereas novel codes capture transient excursions. Under this framing, high reconstruction with poor temporal structure in standard SAEs reflects a prior that fragments local manifold geometry (as formalized in Sec. 4); adding the temporal prior helps preserve the geometry that yields temporally coherent representations (Berkes and Wiskott, 2005; Olshausen and Cadieu, 2007).
@@
Across narrative and dialogue settings, the Temporal Feature Analyzer’s predictable code organizes representations into hierarchical blocks and smooth paths—an effect reminiscent of “temporal straightening”  (Hénaff et al., 2019; Hosseini and Fedorenko, 2023; Xu et al., 2025), where the representational trajectory of an input sequence becomes linearized to support extrapolation toward future states of the sequence. In stories, predictive codes align with (sub)event boundaries and show robustness under noise; in OOD dialogue, they cleanly separate roles and produce near-geodesic single-conversation paths. Kernel spectra and CKA further support the slow/fast split: predictive codes align with slow components; novel codes and standard SAEs align with fast ones. We believe this dynamic warrants further investigation, and may serve fertile ground to develop computational hypotheses about temporal straightening itself.
@@ Acknowledgments
ESL thanks members of the Mechanisms team at Goodfire AI, particularly Tom McGrath, Owen Lewis, Jack Merullo, and Atticus Geiger for helpful discussions. ESL additionally thanks Jack Lindsey for useful comments that helped formalize the intrinsic dimensionality arguments in Sec. 4. The authors also thank David Bau for a helpful discussion about the generalization ability of SAEs. ESL and SSRH further thank Noor Sajid and members of the CRISP lab at Harvard for helpful conversations. CR is supported by a MATS extension grant. AM’s work is partially supported by the National Science Foundation (Grant No. 2530728) and Binational Science Foundation. GT’s work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute at Harvard University.
@@ A.2 Analysis of Stories: UMAP Projections, Dendrograms, Similarity Heatmaps, and Event Boundary Detection
For experiments in Sec. 6.1, we used either stories sampled from the TinyStories dataset [Eldan and Li, 2023] (for Fig. 56) or synthetically sampled stories using GPT-5 (for Fig. 7). For the former, we sampled stories that were up to a 100 tokens long, allowing us to characterize compute UMAP for both Gemma and Llama models without hitting memory bottlenecks. For consistency, we then used these stories for all experiments. For synthetically defined stories, we defined 3 in-context examples (one of which is shown in Fig. 7a) and prompted GPT-5 to produce 50 stories with similar such suddenly changing events. Stories varied in size, but were generally less than 150 tokens long. Following Georgiou et al. [2023], the generated stories’ events were labeled using GPT-5. These labels served as the ‘ground truth’ event boundaries. We qualitatively analyzed and confirmed the event boundaries overlap with our intuitively expected event structure in the stories.
@@
To isolate the low-dimensional geometry of latent codes, we perform a 3D UMAP analysis using the open-source package [McInnes and Healy, 2025]. Similarly, Dendrograms in Fig. 5 are computed using the hierarchical clustering package in SciPy SciPy [2025]. Branches are colored based on proximity, which in our case was between 011 (since we use cosine similarity as the clustering measure). We put a proximity bound of $0.2$ as the distance under which branches are colored within a cluster. For Fourier Analysis, as mentioned in Sec. 6.1, we performed a Fourier transform of the model activations for a given story, isolated its lower frequency components—defined as $0.1\times$ the Nyquist rate, which turns out to possess $\sim$50% of the signal norm—and compute the cosine similarity kernel of the low / high frequency spectra. These kernels are then compared to kernels of latent codes derived using Temporal Feature Analysis or SAEs.
@@ A.3 Analyzing Garden Path Sentences: Dendrograms and Phrase Similarity
For assessing whether tokens in verb phrase relate more with the subject phrase, as would be expect by the typical parse in our used garden path sentences, or with the object phrase, as would be necessary for the correct parse, we computed Dendrograms using the hierarchical clustering package in SciPy [SciPy, 2025]. The similarity between subject phrase (SP), verb phrase (V), and object phrase (OP) was computed by taking the set of tokens $T_{P}$ that belong to a phrase $P\in\{$SP, V, OP}\}, and computing the average latent code: $\bar{c}(P)=\frac{1}{|P|}\sum_{t\in T_{P}}c_{t}$, where $c_{t}$ denotes the latent code extracted for token $t$ using either SAEs or the predictive or novel component of Temporal Feature Analysis. This protocol is similar to popularly used strategies for computing sentence embeddings (see, e.g., work using BERT embeddings [Koroteev, 2021]). Cosine similarity between these phrase-averaged garden path / control sentences yields the tables shown in Fig. 8.
@@ A.4 Analyzing In-Context Representations: Visualization and Counterfactual Experiments
We followed the pipeline of Park et al. [2025] and defined grid graphs where random 1-token words are associated to the nodes of a $\times3$ grid. A random walk of length 500 tokens is performed on the grid, with the visited nodes stringed together with a whitespace (e.g., stringing apple, car, house yields the string ‘apple car house’). These walks were inputted into our analyzed, pretrained language models (Gemma-2-2B in Fig. 9 and Llama-3.1-8B in Fig. 52) to extract corresponding activations.
@@ B.1 Training, Harvesting, and Inference for Temporal Feature Analysis and SAEs
All analyzed SAEs were trained from scratch on 1B precached activations from the Pile-Uncopyrighted dataset [Monology, 2021]. While we did try to use existing, off-the-shelf SAEs, we found the results to be wildly inconsistent depending on where we borrowed the SAE from. To enable a consistent and fair evaluation, we thus preferred to train all SAEs from scratch. For Gemma models, activations were extracted from Layer 12; for Llama models, from layer 15 (all 0-indexed).
@@
Following best practices, we normalize activations to unit expected norm [Bricken et al., 2023, Costa et al., 2025]. This helps put different SAEs on the same training scale, and especially for ReLU SAEs, makes training substantially easier by reducing the need for tuning regularization strength. On a related note, we emphasize our ReLU SAEs have a slightly higher $L_{0}$, i.e., are less sparse, than almost all other SAEs analyzed in this work.
B.2 General Discussion on Training Dynamics and Possible Variants
Since the goal of proposing Temporal Feature Analysis was to identify and highlight the benefits of accommodating the temporal structure of model activations, we did not perform significant tuning of the proposed protocol. That is, we strongly emphasize that our proposed protocol is just an example and we hope much better variants will likely follow as the community explores this research direction—as, e.g., happened for SAEs, which started with vanilla ReLU variants [Cunningham et al., 2023, Bricken et al., 2023], leading to much more sophisticated approaches eventually [Rajamanoharan et al., 2024, Bussmann et al., 2024, Costa et al., 2025]. To assist this exploration, we discuss the following three ideas.
  • Competition dynamic in the learning process. The precise way the training objective is defined in the current version of Temporal Feature Analysis (see Eq. 4), there is a competition dynamic imposed between the predictive and novel component for explaining model representations. Specifically, we found that learning dynamics underwent three phases. First, since the objective of predicting a future direction is more challenging than reconstruction to begin with, we found that the novel part, which is defined using a standard SAE and can in fact reconstruct a bulk of the training signal (since SAEs are trainable in the first place), will dominate the learning dynamics. Then, once the dictionary quality improves, the ability to predict starts to improve and the learning dynamics goes through a chaotic phase with successful training resulting in the predictive component taking over the learning dynamics. Finally, we saw that the novel component keeps conceding its dominance over the reconstructed signal, such that its NMSE finally becomes very high and the predictive component’s NMSE becomes very low. Unsuccessful training, however, can also happen here: at times we saw a fourth phase occurs where the novel component, relatively suddenly, takes over the learning dynamics. This would make the predictive component futile. This was rarely seen when working with Gemma models, but a frequent occurrence in Llama models; we resolved this problem by switching from using TopK to BatchTopk activation for the novel component, which resulted in perfectly stable training for Llama. We believe there are principled mechanisms to address this challenge, however. For example, Temporal Feature Analysis, in its structure, is very similar to Kalman filters [Welch et al., 1995]—latter generally involves Markov assumption, but Temporal Feature Analysis uses the entire context. Seeking inspiration from training objectives for Kalman filters, it can be useful to decompose the loss into two halves: (i) a term that uses only the predictive component to reconstruct the signal, and (ii) a term that uses the novel component to reconstruct just the residual. This breaks the competition dynamic by introducing an asymmetry in the training objective. Another alternative can be the use of StopGrad, as was done in self-supervised learning models of vision that observed similar challenges in training dynamics [Chen and He, 2021, Caron et al., 2021].
  • Rank sparsity. While the predictive component is dense in the sense used in interpretability literature, i.e., its $L_{0}$ norm is not constrained and can be high, in practice we found the learned attention maps and cosine similarity matrices (e.g., in Sec. 6.1) to be very low-rank (e.g., see Fig. 12). In fact, we note in the broader field of dictionary learning and sparse coding [Elad, 2010], an alternative notion of sparsity is rank-sparsity. Specifically, in this paradigm, one constrains the rank of a matrix to be low: in the Euclidean sense, the latent code to reconstruct the target signal will still be dense, but the number of directions used to achieve that, in an alternative space, will be small. While one can optimize for this property using a nuclear norm like regularization objective, this can be expensive in practice (due to the need of an SVD). An alternative is thus to constrain the dimensions of the factors used for a signal’s reconstruction, i.e., what is often called imposing a “rank bottleneck”. When using this methodology on the predictive component, i.e., reducing the dimensions in the K, Q, V matrices of the attention layer, we find the resulting Temporal Feature Analyzers to achieve similar reconstruction as the unconstrained ones. While we did not explore this thread further, we believe it can be promising and is related to recent work on interpreting attention layers [He et al., 2025, Kamath et al., 2025].
  • Layer choice. For all our experiments, we trained SAEs and Temporal Feature Analyzers at layers found around 50% model depth (Layer 12 for Gemma and Layer 15 for Llama; both 0-indexed). When we tried to train Temporal Feature Analyzers on a deeper layer—specifically, layer 26 of Llama—we found the predictive component would not be dominate the norm of the latent code, contributing around 40%. We believe this is a feature, not a bug. In particular, deeper layers have to unembed the extracted contextual signal to make a next-token prediction, and correspondingly a model is likely to discard unnecessary information in the deeper layer that can interfere with its ability to confidently produce next-token predictions. This claim is corroborated by block structured similarities seen in CKA and other representation similarity analysis literature [Kornblith et al., 2019, Raghu et al., 2017]. We believe this thread warrants further investigation, as it may hint at a deeper result on temporal dynamics be varied across different layers, such that one could argue SAEs have the right inductive biases for interpreting deeper layers, but not middle or earlier layers.
Figure 12: Attention Patterns. Attention weights for single sequences across three domains.
@@ H.0.1 ReLU SAE
The vanilla ReLU SAE (Bricken et al. [2023], Cunningham et al. [2023]) is trained with the $L_{1}$-norm penalty:

### v2

@@ 8 Discussion
A longstanding view in computational and cognitive neuroscience treats perception as predictive and time-coupled: slowly varying, context-invariant structure coexists with fast-changing, surprising residuals (levy2008expectation; millidge2024predictive; shain2020fmri). We find that this structure is reflected in LM activations (Sec. 3): activations are strongly correlated with recent context and can be decomposed accordingly. In contrast, SAEs assume independence over time; this is misaligned with the identified non-stationary structure in language model activations. Our protocol—Temporal Feature Analysis—relaxes this assumption, incorporating empirically observed temporal correlations to surface structure that i.i.d. priors obscure. We showed that Temporal Feature Analysis is able to highlight temporal structures in its predictable codes that SAEs are unable to capture due to their prior-data mismatch. Overall, our work adds to recent findings that inductive biases about the specific application under study are necessary to build improved interpretability tools (hindupur2025projecting; costa2025flat; sutter2025the).
@@ 8.1 Future Work
We believe our work brings interpretability of language models one step closer to recent advances in interpretability of neural networks designed to simulate computational neuroscience tasks (sussillo2013opening; cotler2023analyzing; mante2013context; maheswaranathan2019universality; maheswaranathan2020recurrent). Specifically, rather than isolated directions, as is assumed in both sparse coding (olshausen1996emergence; olshausen1997sparse) and SAEs (elhage2022superposition; bricken2023monosemanticity), “features” are better viewed as low-dimensional manifolds that evolve over time—as has also been stated in other recent work (modell2025origins; fel2025into; gurnee2025when). In particular, we see in our results that predictive codes trace smooth, stable trajectories (events, roles, long-horizon constraints), whereas novel codes capture transient excursions. Under this framing, high reconstruction with poor temporal structure in standard SAEs reflects a prior that fragments local manifold geometry (as formalized in Sec. 4); adding the temporal prior helps preserve the geometry that yields temporally coherent representations (berkes2005slow; olshausen2007learning).
@@
Across narrative and dialogue settings, the Temporal Feature Analyzer’s predictable code organizes representations into hierarchical blocks and smooth paths—an effect reminiscent of “temporal straightening”  (henaff2019perceptual; hosseini2023large; xutemporal), where the representational trajectory of an input sequence becomes linearized to support extrapolation toward future states of the sequence. In stories, predictive codes align with (sub)event boundaries and show robustness under noise; in OOD dialogue, they cleanly separate roles and produce near-geodesic single-conversation paths. Kernel spectra and CKA further support the slow/fast split: predictive codes align with slow components; novel codes and standard SAEs align with fast ones. We believe this dynamic warrants further investigation, and may serve fertile ground to develop computational hypotheses about temporal straightening itself.
@@
Training Dynamics and Possible Variants of Temporal Feature Analysis
Since the goal of proposing Temporal Feature Analysis was to identify and highlight the benefits of accommodating the temporal structure of model activations, we did not perform significant tuning of the proposed protocol. That is, we strongly emphasize that our proposed protocol is just an example and we hope much better variants will likely follow as the community explores this research direction—as, e.g., happened for SAEs, which started with vanilla ReLU variants (cunningham2023sparse; bricken2023monosemanticity), eventually leading to much more sophisticated approaches eventually (rajamanoharan2024jumpingaheadimprovingreconstruction; bussmann2024batchtopksparseautoencoders; costa2025flat). To assist this, we highlight the following.
Figure 12: Competition Dynamics. We see while the overall learning curve looks smooth, the predictive and novel code in fact compete with each other to explain the input signal.
1. Competition dynamic in the learning process. The precise way the training objective is defined in the current version of Temporal Feature Analysis (see Eq. 4), there is a competition dynamic imposed between the predictive and novel component for explaining model representations. Specifically, we found that learning dynamics underwent three phases (see Fig. 12). First, since the objective of predicting a future direction is more challenging than reconstruction to begin with, we found that the novel part, which is defined using a standard SAE and can in fact reconstruct a bulk of the training signal (since SAEs are trainable in the first place), will dominate the learning dynamics. Then, once the dictionary quality improves, the ability to predict starts to improve and the learning dynamics goes through a chaotic phase with successful training resulting in the predictive component taking over the learning dynamics. Finally, we saw that the novel component keeps conceding its dominance over the reconstructed signal, such that its NMSE finally becomes very high and the predictive component’s NMSE becomes very low. Unsuccessful training, however, can also happen here: at times we saw a fourth phase occurs where the novel component, relatively suddenly, takes over the learning dynamics. This would make the predictive component futile. This was rarely seen when working with Gemma models, but a frequent occurrence in Llama models; we resolved this problem by switching from using TopK to BatchTopk activation for the novel component, which resulted in perfectly stable training for Llama. We believe there are principled mechanisms to address this challenge, however. For example, Temporal Feature Analysis, in its structure, is very similar to Kalman filters (welch1995introduction)—latter generally involves Markov assumption, but Temporal Feature Analysis uses the entire context. Seeking inspiration from training objectives for Kalman filters, it can be useful to decompose the loss into two halves: (i) a term that uses only the predictive component to reconstruct the signal, and (ii) a term that uses the novel component to reconstruct just the residual. This breaks the competition dynamic by introducing an asymmetry in the training objective.
Figure 13: Attention Patterns. Attention weights for single sequences across three domains show low-rank structure in similarities.
2. Rank sparsity. While the predictive component is dense in the sense used in interpretability literature, i.e., its $L_{0}$ norm is not constrained and can be high, in practice we found the learned attention maps and cosine similarity matrices (e.g., in Sec. 6.1) to be very low-rank (e.g., see Fig. 13). In fact, we note in the broader field of dictionary learning and sparse coding (elad2010sparse), an alternative notion of sparsity is rank-sparsity. Specifically, in this paradigm, one constrains the rank of a matrix to be low: in the Euclidean sense, the latent code to reconstruct the target signal will still be dense, but the number of directions used to achieve that, in an alternative space, will be small. While one can optimize for this property using a nuclear norm like regularization objective, this can be expensive in practice (due to the need of an SVD). An alternative is thus to constrain the dimensions of the factors used for a signal’s reconstruction, i.e., what is often called imposing a “rank bottleneck”. When using this methodology on the predictive component, i.e., reducing the dimensions in the K, Q, V matrices of the attention layer, we find the resulting Temporal Feature Analyzers to achieve similar reconstruction as the unconstrained ones. While we did not explore this thread further, we believe it can be promising and is related to recent work on interpreting attention layers (he2025towards; kamath2025tracing).
Figure 14: Llama Learning Curves. Temporal Feature Analysis works best in middle layers of the model, with training on deeper layers often showing a learning dynamic whereby the novel part eventually takes over and starts to explain a bulk of the signal.
3. Layer choice. For all our experiments, we trained SAEs and Temporal Feature Analyzers at layers found around 50% model depth (Layer 12 for Gemma and Layer 15 for Llama; both 0-indexed). When we tried to train Temporal Feature Analyzers on a deeper layer—specifically, layer 26 ($\sim$81% depth) of Llama—we found the predictive component would not be dominate the norm of the latent code (see Fig. 14). We believe this is a feature, not a bug. In particular, deeper layers have to unembed the extracted contextual signal to make a next-token prediction, and correspondingly a model is likely to discard unnecessary information in the deeper layer that can interfere with its ability to confidently produce next-token predictions. This claim is corroborated by block structured similarities seen in CKA and other representation similarity analysis literature (kornblith2019similarity; raghu2017svcca). We believe this thread warrants further investigation, as it may hint at a deeper result on temporal dynamics be varied across different layers, such that one could argue SAEs have the right inductive biases for interpreting deeper layers, but not middle or earlier layers.
@@ Acknowledgments
ESL thanks members of the Mechanisms team at Goodfire AI, particularly Tom McGrath, Owen Lewis, Jack Merullo, and Atticus Geiger for helpful discussions. ESL additionally thanks Jack Lindsey for useful comments that helped formalize the intrinsic dimensionality arguments in Sec. 4. The authors also thank David Bau for a helpful discussion about the generalization ability of SAEs. ESL and SSRH further thank Noor Sajid and members of the CRISP lab at Harvard for helpful conversations. CR is supported by a MATS extension grant. AM’s work is partially supported by the National Science Foundation (Grant No. 2530728) and Binational Science Foundation. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute at Harvard University.
@@ A.2 Analysis of Stories: UMAP Projections, Dendrograms, Similarity Heatmaps, and Event Boundary Detection
For experiments in Sec. 6.1, we used either stories sampled from the TinyStories dataset [eldan2023tinystories] (for Fig. 56) or synthetically sampled stories using GPT-5 (for Fig. 7). For the former, we sampled stories that were up to a 100 tokens long, allowing us to characterize compute UMAP for both Gemma and Llama models without hitting memory bottlenecks. For consistency, we then used these stories for all experiments. For synthetically defined stories, we defined 3 in-context examples (one of which is shown in Fig. 7a) and prompted GPT-5 to produce 50 stories with similar such suddenly changing events. Stories varied in size, but were generally less than 150 tokens long. Following georgiou2023using, the generated stories’ events were labeled using GPT-5. These labels served as the ‘ground truth’ event boundaries. We qualitatively analyzed and confirmed the event boundaries overlap with our intuitively expected event structure in the stories.
@@
To isolate the low-dimensional geometry of latent codes, we perform a 3D UMAP analysis using the open-source package [umap]. Similarly, Dendrograms in Fig. 5 are computed using the hierarchical clustering package in SciPy dendrogram. Branches are colored based on proximity, which in our case was between 011 (since we use cosine similarity as the clustering measure). We put a proximity bound of $0.2$ as the distance under which branches are colored within a cluster. For Fourier Analysis, as mentioned in Sec. 6.1, we performed a Fourier transform of the model activations for a given story, isolated its lower frequency components—defined as $0.1\times$ the Nyquist rate, which turns out to possess $\sim$50% of the signal norm—and compute the cosine similarity kernel of the low / high frequency spectra. These kernels are then compared to kernels of latent codes derived using Temporal Feature Analysis or SAEs.
@@ A.3 Analyzing Garden Path Sentences: Dendrograms and Phrase Similarity
For assessing whether tokens in verb phrase relate more with the subject phrase, as would be expect by the typical parse in our used garden path sentences, or with the object phrase, as would be necessary for the correct parse, we computed Dendrograms using the hierarchical clustering package in SciPy [dendrogram]. The similarity between subject phrase (SP), verb phrase (V), and object phrase (OP) was computed by taking the set of tokens $T_{P}$ that belong to a phrase $P\in\{$SP, V, OP}\}, and computing the average latent code: $\bar{c}(P)=\frac{1}{|P|}\sum_{t\in T_{P}}c_{t}$, where $c_{t}$ denotes the latent code extracted for token $t$ using either SAEs or the predictive or novel component of Temporal Feature Analysis. This protocol is similar to popularly used strategies for computing sentence embeddings (see, e.g., work using BERT embeddings [koroteev2021bert]). Cosine similarity between these phrase-averaged garden path / control sentences yields the tables shown in Fig. 8.
@@ A.4 Analyzing In-Context Representations: Visualization and Counterfactual Experiments
We followed the pipeline of park2025iclrincontextlearningrepresentations and defined grid graphs where random 1-token words are associated to the nodes of a $\times3$ grid. A random walk of length 500 tokens is performed on the grid, with the visited nodes stringed together with a whitespace (e.g., stringing apple, car, house yields the string ‘apple car house’). These walks were inputted into our analyzed, pretrained language models (Gemma-2-2B in Fig. 9 and Llama-3.1-8B in Fig. 54) to extract corresponding activations.
@@ B.1 Training, Harvesting, and Inference for Temporal Feature Analysis and SAEs
All analyzed SAEs were trained from scratch on 1B precached activations from the Pile-Uncopyrighted dataset [monology2021pile-uncopyrighted]. While we did try to use existing, off-the-shelf SAEs, we found the results to be wildly inconsistent depending on where we borrowed the SAE from. To enable a consistent and fair evaluation, we thus preferred to train all SAEs from scratch. For Gemma models, activations were extracted from Layer 12; for Llama models, from layer 15 (all 0-indexed).
@@
Following best practices, we normalize activations to unit expected norm [bricken2023monosemanticity, costa2025flat]. This helps put different SAEs on the same training scale, and especially for ReLU SAEs, makes training substantially easier by reducing the need for tuning regularization strength. On a related note, we emphasize our ReLU SAEs have a slightly higher $L_{0}$, i.e., are less sparse, than almost all other SAEs analyzed in this work.
@@ H.0.1 ReLU SAE
The vanilla ReLU SAE (bricken2023monosemanticity, cunningham2023sparse) is trained with the $L_{1}$-norm penalty: